ABC Ltd issues a $18 million IPO providing proceeds to ABC of $2.7 per share, from an offer price to the public of $3 per share. The company's legal fees, ASIC registration fees, and other administrative costs are $441,000. The company's share price increases 14 per cent on the first day. What is the underwriting cost? (in millions of dollars to the nearest three decimal places; don't use the $ sign eg 7.897)

When ABC Company went public in September 2008, the offer price was $2.36 per share and the closing price at the end of the first day was $4.86. The company issued 7 million shares. What was the loss to the company due to under-pricing? (in millions of dollars to the nearest two decimal places; don't use $ sign eg $4.5766 million is 4.58)

ABC Ltd, a high-technology company, issues a $34 million IPO with an offer price of $4 per share, underwritten at $3.76 per share. The company's legal fees, ASIC registration fees, and other administrative costs are $394,000. The company's share price increases by $0.6 on the first day. What is the company's total cost of issuing the securities (in millions of dollars to three decimal places; don't use $ sign eg $4.5766 million is 4.577)? (Remember to round the number of shares issued to a whole number)
Answer #1

1]

Underwriting cost = gross issue amount * (offer price per share - net proceeds per share) / offer price per share.

Underwriting cost = $18 million * ($3 - $2.7) / $3

Underwriting cost = $1,800,000, or $1.800 million.

2]

Loss due to underpricing = (closing price - offer price) * shares issued.

Loss due to underpricing = ($4.86 - $2.36) * 7 million = $17.500 million.

3]

total cost = underwriting cost + admin/legal/other costs + underpricing cost.

Number of shares issued = gross issue amount / offer price per share.

Number of shares issued = $34 million / $4 = 8,500,000.

Underwriting cost = shares issued * (offer price - underwritten price)

Underwriting cost = 8,500,000 * ($4 - $3.76) = $2,040,000.

Underpricing cost =  increase in share price * shares issued.

Underpricing cost = $0.6 * 8,500,000 = $5,100,000.

total cost = $2,040,000 + $394,000 + $5,100,000

total cost = $7,534,000, or $7.534 million.